# Hyper-Compression: Model Compression via Hyperfunction

### v3

@@ 3.2 Ergodic Theory
We underscore that designing a suitable hyperfunction is an open-ended question. Generically, the selection of the hyperfunction should balance the compression time, computational complexity, and the compression rate. Unexpectedly, we find the connection between the hyperfunction and ergodic theory to address this question from a unique angle. A branch of ergodic theory Cornfeld et al. (2012) studies on what conditions a low-dimensional dynamic system’s trajectory can cover all points in a high-dimensional space. If a low-dimensional system’s trajectory can cover all points in a high-dimensional space, from an engineering perspective, it is feasible to use low-dimensional curves to approximate high-dimensional points. Thus, we can use fewer parameters to denote more parameters via explicitly characterizing the target point with the low-dimensional dynamic system’s trajectory. Now, we chart the pathway formally:
Definition 1 (Cornfeld et al. (2012)).
@@
Definition 2 (Cornfeld et al. (2012)).
A measure-preserving transformation is a quartet $(X,\mathscr{B},\mu,T)$, where $(X,\mathscr{B},\mu)$ is a measure space, and 1. $T$ is measurable: $E\in\mathscr{B}\Rightarrow T^{-1}E\in\mathscr{B}$; 2. $m$ is $T$invariant: $\mu\left(T^{-1}E\right)=\mu(E)$ for all $E\in\mathscr{B}$.
@@
Given a set $A$ with $\mu(A)>0$, for any target point $\boldsymbol{x}\in\mathbb{R}^{N}$ and $\epsilon>0$, there exists a point $\boldsymbol{x}_{0}\in A$ and $k\in\mathbb{Z}$ such that
$\|T^{k}(\boldsymbol{x}_{0})-\boldsymbol{x}\|<\epsilon.$ (3)
@@
Let two sets $B_{\epsilon/2}(\boldsymbol{x}_{0})\in A$ and $B_{\epsilon}(\boldsymbol{x})\in A$, where $B_{\epsilon}(\boldsymbol{y})=\{\boldsymbol{z}~{}|~{}\|\boldsymbol{z}-\boldsymbol{y}\|<\epsilon\}$. Due to Theorem 1 and the fact that $T$ is measure-preserving, there exists $k$ such that $T^{k}(B_{\epsilon/2}(\boldsymbol{x}_{0}))\in B_{\epsilon}(\boldsymbol{x})$ almost for every point from $B_{\epsilon/2}(\boldsymbol{x}_{0})$, which concludes the proof. ∎
Figure 2: Based on ergodicity, we can construct a deterministic relationship between the one-dimensional trajectory quantity $k$ and the high-dimensional target point $\boldsymbol{x}\in\mathbb{R}^{N}$. Thus, we can compress $\boldsymbol{x}$ into $k$, and decompress $\boldsymbol{x}$ from $k$ based on $T^{k}(\boldsymbol{x}_{0})$.
Let us interpret Theorem 1 and Corollary 2, with an emphasis on how it is related to the model compression problem. First, Definitions 2-3 and Theorem 1 are purely set-theoretic, and hold true for arbitrarily high-dimensional space. This means that Corollary 2 also holds true for an arbitrarily large $N$. Second, Theorem 1 shows that when the measure-preserving transformation is ergodic, even though $A$ is a tiny set, as long as its measure is positive, applying $T$ on $A$ recursively will create a one-dimensional trajectory which can fill the entire space $X$. As a result, per Corollary 2, we can construct a deterministic relationship between the one-dimensional composition number $k$ and the high-dimensional target point $\boldsymbol{x}\in\mathbb{R}^{N}$, as shown in Figure 2. Third, based on Corollary 2, we can compress a group of parameters $\boldsymbol{x}\in\mathbb{R}^{N}$ into a number $k$ by
$\boldsymbol{x}\approx T^{k}(\boldsymbol{x}_{0}),$ (4)
where $\boldsymbol{x}$ is the target. Element-wise, we have
$x_{n}\approx[T^{k}(\boldsymbol{x}_{0})]_{n},n=1,2,\mathinner{\mkern-0.1mu\cdotp\mkern-0.3mu\cdotp\mkern-0.3mu\cdotp\mkern-0.1mu},N.$ (5)
Again, the prerequisite of this kind of compression is $\texttt{memory}(k)<\texttt{memory}(\boldsymbol{x})$.
Theorem 1 and Corollary 2 provide a natural and elegant way to encode a group of numbers into one number. However, we should only be cautiously optimistic about its effect in compression, since simply stitching digits of numbers can also ensemble a group of numbers into one number $(p_{1}p_{2}p_{3},q_{1}q_{2}q_{3})\to p_{1}p_{2}p_{3}q_{1}q_{2}q_{3}$. What matters is whether we can find a short $k$ to approximate $x_{1},\mathinner{\mkern-0.1mu\cdotp\mkern-0.3mu\cdotp\mkern-0.3mu\cdotp\mkern-0.1mu},x_{N}$ based on Eq. (6), such that the memory footprint of the former is smaller than the latter.
Remark 1. The idea of compressing parameters via ergodic theory can also be generalized to continuous dynamic systems which, for example, define the transformation $T$ with differential equations. In this case, the composition number $k$ turns into the trajectory length. In addition, $T$ can be either parametric or non-parametric. Since $T$ is universal for all the given $\boldsymbol{x}$, parameters defining $T$ only have a moderate impact on the compression rate but may provide more flexibility in compressing parameters.
@@ 3.3 Specific Case
Earlier, we outlined a generic framework for the hyper-compression via ergodic theory. The next step is to examine the specific methodologies employed within this framework. Besides the basic ergodicity, the most important question is which dynamic system fits this specific compression issue most? By addressing this question, we can translate the theory to compress the network parameter vector $\boldsymbol{w}$.
@@
Theorem 3 (Katok et al. (1995)).
Suppose $a_{1},\mathinner{\mkern-0.1mu\cdotp\mkern-0.3mu\cdotp\mkern-0.3mu\cdotp\mkern-0.1mu},a_{N}$ are irrationally independent, for any given set $\{w_{n}\}_{n=1}^{N}\subseteq[0,1]$ and $\epsilon>0$, there exists a value $\theta^{*}\in[0,+\infty)$ such that
$|w_{n}-\tau(\theta^{*}a_{n})|<\epsilon,\;n=1,\mathinner{\mkern-0.1mu\cdotp\mkern-0.3mu\cdotp\mkern-0.3mu\cdotp\mkern-0.1mu},N,$ (6)
@@
Figure 3: A two-dimensional example to explain Theorem 3 that one-dimensional continuous dynamic system can fill the high-dimensional space. The irrational independence ensures that the trajectory never overlap. Thus, given a two-dimensional point $[\boldsymbol{w}_{1},\boldsymbol{w}_{2}]$, we can find the corresponding $\theta$ to approximate $[\boldsymbol{w}_{1},\boldsymbol{w}_{2}]$ by $[\tau(\theta/(\pi+1)),\tau(\theta/(\pi+2))]$. For example, $[0.07405,0.00623]\approx[\tau(108/(\pi+1)),\tau(108/(\pi+2))]$.
It can be seen that Theorem 3 is a special case of Eq. (2): Here, the initial point $\boldsymbol{x}_{0}$ is the original point. In addition, because Theorem 3 is a continuous dynamic system, the trajectory parameter $\theta$ corresponds to the composition number $k$ of Eq. (2). Theorem 3 is a corollary of Lemma 19 from Zhang et al. (2022a). As shown in Figure 3, the irrational independence ensures that the trajectory never overlaps so that it can eventually fill the high-dimensional space, when $\theta$ goes to the infinity. In this theorem, the irrational independence condition is readily fulfilled; for instance, one can simply select $a_{k}=1/(\pi+k)$. Later, we empirically and theoretically illustrate that adjusting $a_{k}$ is instrumental in reducing the approximation error and minimizing the performance drop.
@@
The entire algorithmic development is anchored in Eq. (6). Mathematically, given a weight vector $\boldsymbol{w}$, we define the hyper-compression via Theorem 3 as
$\mathcal{H}(\boldsymbol{w}):=\arg\min_{\theta\leq\Theta,~{}\theta,\Theta\in\mathbb{Z}}\|\boldsymbol{w}-\tau(\Delta\cdot\theta\cdot\boldsymbol{a})\|$ (7)
@@
$\mathcal{H}^{-1}(\theta):=\tau(\Delta\cdot\theta\cdot\boldsymbol{a}).$ (8)
@@
Lastly, Theorem 3 needs that $a_{1},a_{2},\mathinner{\mkern-0.1mu\cdotp\mkern-0.3mu\cdotp\mkern-0.3mu\cdotp\mkern-0.1mu},a_{N}$ are irrational numbers. However, in computers that are restricted by the precision limit, all numbers are rational. Theoretically, this will cause the trajectory defined in Theorem 3 to fail covering the entire space. But we actually use Theorem 3 to do approximation, machine precision has been much higher than we need to provide a satisfactory approximation for network parameters.
@@ 4 Error Analysis and Algorithmic Design
In this section, we derive the error bound for using Theorem 1 to compress a single-layer network, which strengthens our understanding for the characters of this kind of compression. Notably, our result can be extended to the multi-layer networks directly. Therefore, we do not put the error bound for these networks for conciseness. Our theory is largely based on Zhang et al. (2023). Hence, we inherit notations from Zhang et al. (2023) for simplicity. Next, we attempt to materialize the idea of using Theorem 1 to do model compression. Drawing insights from our theoretical analysis, we propose four engineering twists to prototype a pragmatic model compression algorithm.
@@ 4.1 Error Analysis
In the single-layer network, we define the input data $X\in\mathbb{R}^{m\times N_{0}}$, the weight matrix $\boldsymbol{W}\in\mathbb{R}^{N_{0}\times N_{1}}$, and the activation function $\phi:\mathbb{R}\to\mathbb{R}$, which means there are $m$ samples, each with $N_{0}$ dimensions, and $N_{1}$ neurons in this single-layer network. Following the divide-and-conquer strategy, estimating the error bound of a single neuron can naturally lead to a good estimation for the entire single network. Without loss of generality, let $\boldsymbol{w}=[w_{1},w_{2},\mathinner{\mkern-0.1mu\cdotp\mkern-0.3mu\cdotp\mkern-0.3mu\cdotp\mkern-0.1mu},w_{N_{0}}]$ be the weight vector of a neuron. With Eqs. (7) and (8), the compression and decompression can be expressed as
$\left\{\begin{array}[]{l}\theta=\mathcal{H}\left(\boldsymbol{w}\right)\\ \boldsymbol{q}=\mathcal{H}^{-1}(\theta),\\ \end{array}\right.$ (9)
$\boldsymbol{q}$ is the decompressed weight vector.
Since the hyper-compression is not lossless, there must be a discrepancy between $\boldsymbol{w}$ and $\boldsymbol{q}$, which results in an error in the neuron’s final output. To characterize this error, we first characterize the pre-activation error of a neuron. Specifically, we denote pre-activation error recursively in the form of error accumulation.
$\left\{\begin{array}[]{l}\boldsymbol{u}_{0}=\mathbf{0}\\ \boldsymbol{u}_{t}=\boldsymbol{u}_{t-1}+w_{t}\boldsymbol{X}_{t}-q_{t}\boldsymbol{X}_{t}\\ \boldsymbol{u}_{N_{0}}=\boldsymbol{X}\boldsymbol{w}-\boldsymbol{X}\boldsymbol{q},\end{array}\right.$ (10)
where $\boldsymbol{u}_{t}$ is the accumulated error generated from $1^{st}$ to the $t$-th weight in the neuron. Thus, $\boldsymbol{u}_{N_{0}}=\boldsymbol{X}\boldsymbol{w}-\boldsymbol{X}\boldsymbol{q}$ is the total error. Our goal is to describe how $\boldsymbol{u}_{N_{0}}$ is bounded. The error is measured by the $L_{1}$-norm, i.e., $\left\|\boldsymbol{u}\right\|_{1}=\left\|\boldsymbol{u}\right\|=\sum_{i}|u_{i}|$.
@@
Given the input data $X\in[0,1]^{m\times N_{0}}$, suppose that each element of the column $X_{t}$ of $X\in\mathbb{R}^{N_{0}}$ is i.i.d. drawn from a uniform distribution over $[0,1]$, and $\boldsymbol{w}\in[0,1]^{N_{0}}$, we compress and decompress $\boldsymbol{w}$ based on Eq. (9) with the step size $\Delta>0$ and the maximum integer $\Theta$. Then, there exists $c\in(1/2,1]$, for any $\epsilon<1$, we have $\boldsymbol{q}$
$\mathrm{P}\left(\|\boldsymbol{X}\boldsymbol{w}-\boldsymbol{X}\boldsymbol{q}\|\leq 2cmN_{0}\epsilon\right)\geq 1-e^{-cmN_{0}\epsilon},$ (11)
where $\boldsymbol{q}$ is the recovered weight vector. Furthermore, if the activation function $\varphi:\mathbb{R}\rightarrow\mathbb{R}$ is $\xi$-Lipschitz continuous, that is, $\mid\varphi(x)-\varphi(y)|\leq\xi|x-y\mid$ for all $x,y\in\mathbb{R}$, then we have
$\mathrm{P}\left(\|\varphi(\boldsymbol{X}\boldsymbol{w})-\varphi(\boldsymbol{X}\boldsymbol{q})\|\leq 2cmN_{0}\epsilon\xi\right)\geq 1-e^{-cmN_{0}\epsilon}$ (12)
@@
$\boldsymbol{u}_{t}$ is composed of $w_{t}-q_{t}$ and $\boldsymbol{X}$. We first estimate them, respectively, and then we provide the estimation for $\boldsymbol{u}_{N_{0}}$ by integrating them.
i) Suppose that $\boldsymbol{w}\in[0,1]^{N_{0}}$, we compress $\boldsymbol{w}$ based on Eq. (7) with the step size $\Delta>0$ and the largest integer $\Theta$. Then, there must exist a sufficiently large $\Theta$ such that for $t=1,2,\ldots,N_{0}$,
@@
This is because when $\Theta$ is large, there will be one element $\{\tau(\Delta\cdot\theta\cdot\boldsymbol{a})\}_{\theta=1}^{\Theta}$ sufficiently close to $\boldsymbol{w}$ due to the density of trajectory.
ii) Let $U$ denote the uniform distribution on $[0,1]$ with Suppose that the random vector $\boldsymbol{X}\in\mathbb{R}^{m}$ with each element randomly drawn from $U$. Then as $m\to\infty$, we have
@@
where $c\in(1/2,1]$. The $l_{1}$-norm of the vector $\boldsymbol{X}$ is $\|\boldsymbol{X}\|_{1}=\sum_{i=1}^{n}X_{i}$, where each $X_{i}$ is i.i.d. uniform on $[0,1]$ Because each element $X_{i}$ is independent from each other, the expected value of $\|\boldsymbol{X}\|$ is
$\mathbb{E}\left[\|\boldsymbol{X}\|\right]=m\cdot\mathbb{E}\left[X_{i}\right]=m/2.$ (15)
@@
$\mathrm{P}\left(\left\|\boldsymbol{u}_{N_{0}}\right\|\geq\alpha\right)$ $=\mathrm{P}\left(e^{\eta\left\|\boldsymbol{u}_{N_{0}}\right\|}\geq e^{\alpha}\right)\leq e^{-\alpha}\mathbb{E}e^{\left\|\boldsymbol{u}_{N_{0}}\right\|}$ (16)
$=e^{-\alpha}\mathbb{E}e^{\left\|\sum_{t=1}^{N_{0}}(\boldsymbol{u}_{t}-\boldsymbol{u}_{t-1})\right\|}$
$\leq e^{-\alpha}\sum_{t=1}^{N_{0}}\mathbb{E}e^{\left\|\boldsymbol{u}_{t}-\boldsymbol{u}_{t-1}\right\|}$
@@
$\mathrm{P}\left(\left\|\boldsymbol{u}_{N_{0}}\right\|\geq 2cmN_{0}\epsilon\right)\leq e^{-cmN_{0}\epsilon},$ (17)
which proves Eq. (11). Then, considering the Lipschitz continuity of the activation function $\varphi$, we have $\|\varphi(\boldsymbol{X}\boldsymbol{w})-\varphi(\boldsymbol{X}\boldsymbol{q})\|\leq\xi\|\boldsymbol{X}\boldsymbol{w}-\boldsymbol{X}\boldsymbol{q}\|$. Integrating it with Eq. (11) proves Eq. (12). ∎
@@
Eq. (12) characterize that the sample size $m$, the dimensionality $N_{0}$ and $\Theta$ dominates the reconstruction error $\varphi(\boldsymbol{X}\boldsymbol{w})-\varphi(\boldsymbol{X}\boldsymbol{q})$. It is straightforward to understand that when the sample size $m$ and the dimensionality $N_{0}$ go higher, more error will be accumulated. In addition, increasing $\Theta$ can add more points into the set $\{\tau(\Delta\cdot\theta\cdot\boldsymbol{a}):\theta\leq\Theta\}$, which will naturally increase the likelihood of finding a point closer to the given $\boldsymbol{w}$.
Remark 3. When no post-hoc retraining is performed, hyper-compression can approximately achieve the same level of error but a higher compression rate compared to quantization. Let us take the Int2 quantization as an example to illustrate this point. As Figure 4 shows, we sample five points (A, B, C, D, E) from the trajectory of the dynamic equation. Because of the irrational direction, the projection of all these five points along both $x$ and $y$ axes will not overlap. Projections of these five points will divide both $x$ and $y$ axes into six segments, e.g., $\Xi_{1x},\Xi_{2x},\mathinner{\mkern-0.1mu\cdotp\mkern-0.3mu\cdotp\mkern-0.3mu\cdotp\mkern-0.1mu},\Xi_{6x}$ and $\Xi_{1y},\Xi_{2y},\mathinner{\mkern-0.1mu\cdotp\mkern-0.3mu\cdotp\mkern-0.3mu\cdotp\mkern-0.1mu},\Xi_{6y}$. We reasonably assume that each segment is at the same level of $1/(2^{\texttt{Int2}}+1)$. If we adjust the irrational direction to make points relatively evenly distributed in the 2D space, the error in approximating most 2D points at positions $x$ is usually no more than one segment, and the error will never exceed a summation of two consecutive segments. The hyper-compression achieves the same level of error as Int2. But the hyper-compression has a double compression rate, since it turns two numbers into an integer from $\{1,2,3,4,5\}$.

### v4

@@ 3.2 Ergodic Theory
We underscore that designing a suitable hyperfunction is an open-ended question. Generically, the selection of the hyperfunction should balance the compression time, computational complexity, and the compression rate. Unexpectedly, we find the connection between the hyperfunction and ergodic theory to address this question from a unique angle. A branch of ergodic theory [12] studies on what conditions a low-dimensional dynamic system’s trajectory can cover all points in a high-dimensional space. If a low-dimensional system’s trajectory can cover all points in a high-dimensional space, from an engineering perspective, it is feasible to use low-dimensional curves to approximate high-dimensional points. Thus, we can use fewer parameters to denote more parameters via explicitly characterizing the target point with the low-dimensional dynamic system’s trajectory. Now, we chart the pathway formally:
Definition 1 ([12]).
@@
Definition 2 ([12]).
A measure-preserving transformation is a quartet $(X,\mathscr{B},\mu,T)$, where $(X,\mathscr{B},\mu)$ is a measure space, and 1. $T$ is measurable: $E\in\mathscr{B}\Rightarrow T^{-1}E\in\mathscr{B}$; 2. $m$ is $T$ -invariant: $\mu\left(T^{-1}E\right)=\mu(E)$ for all $E\in\mathscr{B}$.
@@
Figure 2: Based on ergodicity, we can construct a deterministic relationship between the one-dimensional trajectory quantity $k$ and the high-dimensional target point $\bm{x}\in\mathbb{R}^{N}$. Thus, we can compress $\bm{x}$ into $k$, and decompress $\bm{x}$ from $k$ based on $T^{k}(\bm{x}_{0})$.
@@
Suppose the $\texttt{supp}(\mu)$ is of full measure, then given a set $A$ with $\mu(A)>0$, for any target point $\bm{x}\in\mathbb{R}^{N}$ and $\epsilon>0$, there exists a point $\bm{x}_{0}\in A$ and $k\in\mathbb{Z}$ such that
$\|T^{k}(\bm{x}_{0})-\bm{x}\|<\epsilon.$ (3)
@@
Let two sets $B_{\epsilon/2}(\bm{x}_{0})\in A$ and $B_{\epsilon}(\bm{x})\in A$, where $B_{\epsilon}(\bm{y})=\{\bm{z}~|~\|\bm{z}-\bm{y}\|<\epsilon\}$. Due to Theorem 1 and the fact that $T$ is measure-preserving, there exists $k$ such that $T^{k}(B_{\epsilon/2}(\bm{x}_{0}))\in B_{\epsilon}(\bm{x})$ almost for every point from $B_{\epsilon/2}(\bm{x}_{0})$, which concludes the proof. ∎
Let us interpret Theorem 1 and Corollary 2, with an emphasis on how it is related to the model compression problem. First, Definitions 2-3 and Theorem 1 are purely set-theoretic, and hold true for arbitrarily high-dimensional space. This means that Corollary 2 also holds true for an arbitrarily large $N$. Second, Theorem 1 shows that when the measure-preserving transformation is ergodic, even though $A$ is a tiny set, as long as its measure is positive, applying $T$ on $A$ recursively will create a one-dimensional trajectory which can fill the entire space $X$. As a result, per Corollary 2, we can construct a deterministic relationship between the one-dimensional composition number $k$ and the high-dimensional target point $\bm{x}\in\mathbb{R}^{N}$, as shown in Figure 2. Third, based on Corollary 2, we can compress a group of parameters $\bm{x}\in\mathbb{R}^{N}$ into a number $k$ by
$\bm{x}\approx T^{k}(\bm{x}_{0}),$ (4)
where $\bm{x}$ is the target. Element-wise, we have
$x_{n}\approx[T^{k}(\bm{x}_{0})]_{n},n=1,2,\cdots,N.$ (5)
Again, the prerequisite of this kind of compression is $\texttt{memory}(k)<\texttt{memory}(\bm{x})$.
Theorem 1 and Corollary 2 provide a natural and elegant way to encode a group of numbers into one number. However, we should only be cautiously optimistic about its effect in compression, since simply stitching digits of numbers can also ensemble a group of numbers into one number $(p_{1}p_{2}p_{3},q_{1}q_{2}q_{3})\to p_{1}p_{2}p_{3}q_{1}q_{2}q_{3}$. What matters is whether we can find a short $k$ to approximate $x_{1},\cdots,x_{N}$ based on Eq. (6), such that the memory footprint of the former is smaller than the latter.
Remark 1. The idea of compressing parameters via ergodic theory can also be generalized to continuous dynamic systems which, for example, define the transformation $T$ with differential equations. In this case, the composition number $k$ turns into the trajectory length. In addition, $T$ can be either parametric or non-parametric. Since $T$ is universal for all the given $\bm{x}$, parameters defining $T$ only have a moderate impact on the compression rate but may provide more flexibility in compressing parameters.
@@ 3.3 Specific Case
Earlier, we outlined a generic framework for the Hyper-Compression via ergodic theory. The next step is to examine the specific methodologies employed within this framework. Besides the basic ergodicity, the most important question is which dynamic system fits this specific compression issue most? By addressing this question, we can translate the theory to compress the network parameter vector $\bm{w}$.
@@
Theorem 3 ([29]).
Suppose $a_{1},\cdots,a_{N}$ are irrationally independent, for any given set $\{w_{n}\}_{n=1}^{N}\subseteq[0,1]$ and $\epsilon>0$, there exists a value $\theta^{*}\in[0,+\infty)$ such that
$|w_{n}-\tau(\theta^{*}a_{n})|<\epsilon,\;n=1,\cdots,N,$ (6)
@@
Figure 3: A two-dimensional example to explain Theorem 3 that one-dimensional continuous dynamic system can fill the high-dimensional space. The irrational independence ensures that the trajectory never overlap. Thus, given a two-dimensional point $\bm{w}_{1},\bm{w}_{2}]$, we can find the corresponding $\theta$ to approximate $\bm{w}_{1},\bm{w}_{2}]$ by $\tau(\theta/(\pi+1)),\tau(\theta/(\pi+2))]$. For example, $0.07405,0.00623]\approx[\tau(108/(\pi+1)),\tau(108/(\pi+2))]$.
It can be seen that Theorem 3 is a special case of Eq. (2): Here, the initial point $\bm{x}_{0}$ is the original point. In addition, because Theorem 3 is a continuous dynamic system, the trajectory parameter $\theta$ corresponds to the composition number $k$ of Eq. (2). Theorem 3 is a corollary of Lemma 19 from [58]. As shown in Figure 3, the irrational independence ensures that the trajectory never overlaps so that it can eventually fill the high-dimensional space, when $\theta$ goes to the infinity. In this theorem, the irrational independence condition is readily fulfilled; for instance, one can simply select $a_{k}=1/(\pi+k)$. Later, we empirically and theoretically illustrate that adjusting $a_{k}$ is instrumental in reducing the approximation error and minimizing the performance drop.
@@
The entire algorithmic development is anchored in Eq. (6). Mathematically, given a weight vector $\bm{w}$, we define the Hyper-Compression via Theorem 3 as
$\mathcal{H}(\bm{w}):=\arg\min_{\theta\leq\Theta,~\theta,\Theta\in\mathbb{Z}}\|\bm{w}-\tau(\Delta\cdot\theta\cdot\bm{a})\|$ (7)
@@
$\mathcal{H}^{-1}(\theta):=\tau(\Delta\cdot\theta\cdot\bm{a}).$ (8)
@@
Lastly, Theorem 3 needs that $a_{1},a_{2},\cdots,a_{N}$ are irrational numbers. However, in computers that are restricted by the precision limit, all numbers are rational. Theoretically, this will cause the trajectory defined in Theorem 3 to fail covering the entire space. But we actually use Theorem 3 to do approximation, machine precision has been much higher than we need to provide a satisfactory approximation for network parameters.
@@ 4 Error Analysis and Algorithmic Design
In this section, we derive the error bound for using Theorem 1 to compress a single-layer network, which strengthens our understanding for the characters of this kind of compression. Notably, our result can be extended to the multi-layer networks directly. Therefore, we do not put the error bound for these networks for conciseness. Our theory is largely based on [55]. Hence, we inherit notations from [55] for simplicity. Next, we attempt to materialize the idea of using Theorem 1 to do model compression. Drawing insights from our theoretical analysis, we propose four engineering twists to prototype a pragmatic model compression algorithm.
@@ 4.1 Error Analysis
In the single-layer network, we define the input data $X\in\mathbb{R}^{m\times N_{0}}$, the weight matrix $\bm{W}\in\mathbb{R}^{N_{0}\times N_{1}}$, and the activation function $\phi:\mathbb{R}\to\mathbb{R}$, which means there are $m$ samples, each with $N_{0}$ dimensions, and $N_{1}$ neurons in this single-layer network. Following the divide-and-conquer strategy, estimating the error bound of a single neuron can naturally lead to a good estimation for the entire single network. Without loss of generality, let $\bm{w}=[w_{1},w_{2},\cdots,w_{N_{0}}]$ be the weight vector of a neuron. With Eqs. (7) and (8), the compression and decompression can be expressed as
[…]
$\bm{q}$ is the decompressed weight vector.
Since the Hyper-Compression is not lossless, there must be a discrepancy between $\bm{w}$ and $\bm{q}$, which results in an error in the neuron’s final output. To characterize this error, we first characterize the pre-activation error of a neuron. Specifically, we denote pre-activation error recursively in the form of error accumulation.
$\left\{\begin{array}[]{l}\bm{u}_{0}=\mathbf{0}\\ \bm{u}_{t}=\bm{u}_{t-1}+w_{t}\bm{X}_{t}-q_{t}\bm{X}_{t}\\ \bm{u}_{N_{0}}=\bm{X}\bm{w}-\bm{X}\bm{q},\end{array}\right$. (10)
where $\bm{u}_{t}$ is the accumulated error generated from $1^{st}$ to the $t$-th weight in the neuron. Thus, $\bm{u}_{N_{0}}=\bm{X}\bm{w}-\bm{X}\bm{q}$ is the total error. Our goal is to describe how $\bm{u}_{N_{0}}$ is bounded. The error is measured by the $L_{1}$-norm, i.e., $\left\|\bm{u}\right\|_{1}=\left\|\bm{u}\right\|=\sum_{i}|u_{i}|$.
@@
Given the input data $X\in[0,1]^{m\times N_{0}}$, suppose that each element of the column $X_{t}$ of $X\in\mathbb{R}^{N_{0}}$ is i.i.d. drawn from a uniform distribution over $[0,1]$, and $\bm{w}\in[0,1]^{N_{0}}$, we compress and decompress $\bm{w}$ based on Eq. (9) with the step size $\Delta>0$ and the maximum integer $\Theta$. Then, there exists $c\in(1/2,1]$, for any $\epsilon<1$, we have $\bm{q}$
$\mathrm{P}\left(\|\bm{X}\bm{w}-\bm{X}\bm{q}\|\leq 2cmN_{0}\epsilon\right)\geq 1-e^{-cmN_{0}\epsilon},$ (11)
where $\bm{q}$ is the recovered weight vector. Furthermore, if the activation function $\varphi:\mathbb{R}\rightarrow\mathbb{R}$ is $\xi$-Lipschitz continuous, that is, $\mid\varphi(x)-\varphi(y)|\leq\xi|x-y\mid$ for all $x,y\in\mathbb{R}$, then we have
$\mathrm{P}\left(\|\varphi(\bm{X}\bm{w})-\varphi(\bm{X}\bm{q})\|\leq 2cmN_{0}\epsilon\xi\right)\geq 1-e^{-cmN_{0}\epsilon}$ (12)
@@
$\bm{u}_{t}$ is composed of $w_{t}-q_{t}$ and $\bm{X}$. We first estimate them, respectively, and then we provide the estimation for $\bm{u}_{N_{0}}$ by integrating them.
i) Suppose that $\bm{w}\in[0,1]^{N_{0}}$, we compress $\bm{w}$ based on Eq. (7) with the step size $\Delta>0$ and the largest integer $\Theta$. Then, there must exist a sufficiently large $\Theta$ such that for $t=1,2,\ldots,N_{0}$,
@@
This is because when $\Theta$ is large, there will be one element $\{\tau(\Delta\cdot\theta\cdot\bm{a})\}_{\theta=1}^{\Theta}$ sufficiently close to $\bm{w}$ due to the density of trajectory.
ii) Let $U$ denote the uniform distribution on $[0,1]$ with Suppose that the random vector $\bm{X}\in\mathbb{R}^{m}$ with each element randomly drawn from $U$. Then as $m\to\infty$, we have
@@
where $c\in(1/2,1]$. The $l_{1}$-norm of the vector $\bm{X}$ is $\|\bm{X}\|_{1}=\sum_{i=1}^{n}X_{i}$, where each $X_{i}$ is i.i.d. uniform on $[0,1]$ . Because each element $X_{i}$ is independent from each other, the expected value of $\|\bm{X}\|$ is
$\mathbb{E}\left[\|\bm{X}\|\right]=m\cdot\mathbb{E}\left[X_{i}\right]=m/2.$ (15)
@@
$\mathrm{P}\left(\left\|\bm{u}_{N_{0}}\right\|\geq\alpha\right)$ $=\mathrm{P}\left(e^{\eta\left\|\bm{u}_{N_{0}}\right\|}\geq e^{\alpha}\right)\leq e^{-\alpha}\mathbb{E}e^{\left\|\bm{u}_{N_{0}}\right\|}$ (16)
$=e^{-\alpha}\mathbb{E}e^{\left\|\sum_{t=1}^{N_{0}}(\bm{u}_{t}-\bm{u}_{t-1})\right\|}$
$\leq e^{-\alpha}\sum_{t=1}^{N_{0}}\mathbb{E}e^{\left\|\bm{u}_{t}-\bm{u}_{t-1}\right\|}$
@@
$\mathrm{P}\left(\left\|\bm{u}_{N_{0}}\right\|\geq 2cmN_{0}\epsilon\right)\leq e^{-cmN_{0}\epsilon},$ (17)
which proves Eq. (11). Then, considering the Lipschitz continuity of the activation function $\varphi$, we have $\|\varphi(\bm{X}\bm{w})-\varphi(\bm{X}\bm{q})\|\leq\xi\|\bm{X}\bm{w}-\bm{X}\bm{q}\|$. Integrating it with Eq. (11) proves Eq. (12).
Eq. (12) characterize that the sample size $m$, the dimensionality $N_{0}$ and $\Theta$ dominates the reconstruction error $\varphi(\bm{X}\bm{w})-\varphi(\bm{X}\bm{q})$. It is straightforward to understand that when the sample size $m$ and the dimensionality $N_{0}$ go higher, more error will be accumulated. In addition, increasing $\Theta$ can add more points into the set $\{\tau(\Delta\cdot\theta\cdot\bm{a}):\theta\leq\Theta\}$, which will naturally increase the likelihood of finding a point closer to the given $\bm{w}$.
Remark 3. When no post-hoc retraining is performed, Hyper-Compression can approximately achieve the same level of error but a higher compression rate compared to quantization. Let us take the Int2 quantization as an example to illustrate this point. As Figure 4 shows, we sample five points (A, B, C, D, E) from the trajectory of the dynamic equation. Because of the irrational direction, the projection of all these five points along both $x$ and $y$ axes will not overlap. Projections of these five points will divide both $x$ and $y$ axes into six segments, e.g., $\Xi_{1x},\Xi_{2x},\cdots,\Xi_{6x}$ and $\Xi_{1y},\Xi_{2y},\cdots,\Xi_{6y}$. We reasonably assume that each segment is at the same level of $1/(2^{\texttt{Int2}}+1)$. If we adjust the irrational direction to make points relatively evenly distributed in the 2D space, the error in approximating most 2D points at positions $x$ is usually no more than one segment, and the error will never exceed a summation of two consecutive segments. The Hyper-Compression achieves the same level of error as Int2. But the Hyper-Compression has a double compression rate, since it turns two numbers into an integer from $\{1,2,3,4,5\}$.